2. The equations of motion of this system are; ma Seats 12Y+7Y+247-6Z-122=0 6Z + 6Z+12Z- 6Y-12Y=f(t) 7W+7W+14W-7Y-14Y
Answer #1

@ the equation of motion of this system ane; 124 + 74 +244 - 62 - 122 = 0 67 + 621 + 122 - 64 - 12Y = f(t) ¥w + 70+ 140Now represent that equation in state variable form. il 124 - 74 + 244 - 62 - 122 - 0 : 12cxi) + F(x0) + 24(X4) 6 (X2) - 126AX + BY R-CX 싸 X) here f Ex x6 Y = input here input = f(t) R output here output - Y = X4. X). (-7/12) (/) o - 2 o75x12 2 O (1